The sum of two numbers is 7, twice their product gives 20. Find the numbers
Found 2 solutions by ikleyn, Alan3354:
Answer by ikleyn(52778)

.

The sum is 7, the product is 10.


The numbers are 2 and 5.


5 seconds to guess.


In Algebra way,


    x*(7-x) = 10

    7x - x^2 = 10

    x^2 - 7x + 10 = 0

    (x-2)*(x-5) = 0,    etc.

Solved.


Answer by Alan3354(69443)

You can think about it for a few seconds and find they're 2 & 5.
===================
Or, try to make it a Master's Thesis:
x*(7 - x) = 10
7x - x^2 = 10
x^2 - 7x + 10 = 0 ----- At this point, you have to think about and see they're 2 & 5. Why add the extra steps? If they are not integers, this would be necessary.
(x-2)*(x-5) = 0
---> 2 & 5
========================
If the numbers are not integers, the steps would be necessary.